OpenStudy (marsmcnugget):

Solve for w. P = 2l + 2w

OpenStudy (danieldjpon3):

What's P?

OpenStudy (marsmcnugget):

thats all the question says it dosnt say what p is

OpenStudy (marsmcnugget):

w=P−2l2w=P−2l2 w=2l−P2w=2l−P2 w=P−lw=P−l w=P−2ll

OpenStudy (marsmcnugget):

thats the answers

OpenStudy (marsmcnugget):

the choices

OpenStudy (danieldjpon3):

k brb

Awolflover1 (awolflover1):

okay do you want to math this out or check your answers?

OpenStudy (marsmcnugget):

i need help solving it i suck at math

Awolflover1 (awolflover1):

alright :)

OpenStudy (marsmcnugget):

but i dont have much time

Awolflover1 (awolflover1):

I'll do it a faster way :) add -21 on both sides \[2l+2w+−2l=p+−2l\]

Awolflover1 (awolflover1):

\[2w=−2l+p\]

Awolflover1 (awolflover1):

now Divide both sides with 2 \[\frac{ 2w }{ 2 }=\frac{ -21+p }{ 2}\]

Awolflover1 (awolflover1):

now w= ?